OpenStudy (anonymous):

What's the value of log-base 6- of the fourth root of 6)??

OpenStudy (rulnick):

= log_6 (6^(1/4)) = (1/4) log_6 6 = (1/4) 1 = 1/4.

OpenStudy (rulnick):

questions?

OpenStudy (anonymous):

no, thanks, it's great!